How many different ways can 4 people be seated in a car with 7 seats?
(Assume that one person has to drive.)
This problem has been frustrating me for the past 10 minutes. It looks so simple yet I can't get it.

Answer by Theo(13342)

my best guess is as follows:

driver's seat has 4 possibilities because there are 4 people and one of them has to be in the driver's seat.

that leaves 6 seats for 3 people.

looking at the first of the 6 seats, there are 3 choices there because there are 3 people left to choose that seat.
looking at the second of the 6 seats, there are 2 choices there because there are 2 people left to choose that seat.
looking at the third seat, there is 1 choice left there because there is one person looking for that seat.

so far you have 4 * 3! possibilities which is equal to 4 * 3 * 2 * 1 = 24.

now if you look at the 6 seats that aren't the driver's seat, only 3 out of the 6 have been chosen.

that, i believe, gets into how many sets of 3 can you make out of 6 where order is not important which is the combination formula.

that formula is 6C3 = (6*5*4) / (1*2*3) which results in 20 possible combinations.

your total number of possibilities is equal to 24 * 20 which would be equal to 480 possible arrangements.

this is extremely difficult to visualize.

i used much smaller examples to confirm the formula was reasonable.

i'll do one of them so you can see the logic i used.

assume 4 seats with 3 people.

number of possible choices for the driver's seat is 3.

that leaves 3 seats to be filled with the 2 remaining people.

the first seat has 2 choices and the second seat has 1 choice.

so far you have 3 * 2 * 1 = 6 possible choices.

since there are 3 seats and only 2 of them can be used at a time, the possible combinations of 2 seats at a time from 3 seats is 3C2 which is equal to 3.

your solution should be 6 * 3 = 18 possible ways the seats can be filled.

this is a big number but not that big that we can't confirm that the formula gave us the correct answer.

i'll do that below:


                         driver seat    seat 1     seat 2     seat 3

people are a, b, c
x is the empty seat


1                             a            b          c          x
2                             a            c          b          x
3                             a            b          x          c
4                             a            c          x          b
5                             a            x          b          c
6                             a            x          c          b


this is the first set with a in the driver's seat.
the second set will have b in the driver's seat.
the third set will have c in the driver's seat.
the total should be 18.


i'll do the second set of 6 so you can see how it works.


7                             b            a          c          x
8                             b            c          a          x
9                             b            a          x          c
10                            b            c          x          a
11                            b            x          a          c
12                            b            x          c          a


another 6 possible combinations will result from placing c in the driver's seat.
the total is 18.
the formula checks out.
i used it with 5 seats and believe it worked there as well unless i made a mistake.

that's why i think it's good.

so we'll do your problem again just to make sure it was calculated correctly.

you have 7 seats and 4 people.
number of possible ways to fill the driver's seat is 4.
number of possible ways to fill the first 3 seats is 3! = 6.
only 3 of the 6 seats are filled.
number of possible ways to fill the 6 seats 3 at a time is 6C3 = 20.
total number of ways is 4 * 6 * 20 = 480.

without doing an exhaustive investigation i'm pretty sure the solutions is good.
if you agree, then that's the way to go.